A railway truck of 100kg mass traveling with a velocity of 7ms-1collides with a second truck of 200kg mass and the two couple automatically and move off together. Calculate the velocity of the coupled trucks and the energy lost to the system; if the second truck has a speed of 5ms-1(a) in the same direction, (b) in the opposite direction and (c) is initially stationary.


1
Expert's answer
2022-03-01T15:18:29-0500

The conservation of momentum law is "\vec{p}_1 + \vec{p_2} = \vec{p}_{\text{tot}}" or "m_1\vec{v}_1 + m_2\vec{v_2} = (m_1+m_2)\vec{v}_{\text{tot}}"

a) In case of the same direction

"m_1{v}_1 + m_2{v_2} = (m_1+m_2){v}_{\text{tot}}, \\
v_{\text{tot}} = \dfrac{m_1{v}_1 + m_2{v_2}}{m_1+m_2} = \dfrac{100\cdot7 + 200\cdot5}{100+200} = 5.67\,\mathrm{m/s}."


The initial total kinetic energy is "0.5m_1{v_1}^2 + 0.5m_2{v_2}^2 = 4950\,\mathrm{J}."

The final kinetic energy is "0.5(m_1+m_2)\cdot v_{\text{tot}}^2 = 4816.7\,\mathrm{J}."

The energy lost is 133 J.


b) In case of the opposite directions

"m_1{v}_1 - m_2{v_2} = (m_1+m_2){v}_{\text{tot}}, \\
v_{\text{tot}} = \dfrac{m_1{v}_1 - m_2{v_2}}{m_1+m_2} = \dfrac{100\cdot7 - 200\cdot5}{100+200} = -1\,\mathrm{m/s}." So the final velocity is directed along the velocity of the second truck.

The final kinetic energy is "0.5(m_1+m_2)\cdot v_{\text{tot}}^2 = 150\,\mathrm{J}."

The energy lost is 4800 J.


c) In case of initially stationary truck

"m_1{v}_1 = (m_1+m_2){v}_{\text{tot}}, \\
v_{\text{tot}} = \dfrac{m_1{v}_1 }{m_1+m_2} = \dfrac{100\cdot7 }{100+200} = 2.33\,\mathrm{m/s}."

The final kinetic energy is "0.5(m_1+m_2)\cdot v_{\text{tot}}^2 = 816.6\,\mathrm{J}."

The energy lost is 4133.3 J.